OpenStudy (zenmo):

Find all solutions of sin2x = ...

OpenStudy (zenmo):

\[\sin2x=\frac{ -\sqrt{3} }{ 2 }\]

OpenStudy (zenmo):

\[2x=\frac{ -\pi }{ 3 }\] \[x=-\frac{ \pi }{ 6 }\]

OpenStudy (jiteshmeghwal9):

If\[sin \theta=sin \alpha\] then \[\theta=n \pi +(-1)^n\alpha\]wher n=0,1,2...

OpenStudy (jiteshmeghwal9):

\[sin2x=sin \frac{-\pi}{3}\]\[2x=n \pi+(-1)^n \frac{-\pi}{3}\]From here can u determine all the slution